# Must Read: A Systematic Survey of Computational Persuasion

### v1

@@ Abstract.
Persuasion is a fundamental aspect of communication, influencing decision-making across diverse contexts, from everyday conversations to high-stakes scenarios such as politics, marketing, and law. The rise of conversational AI systems has significantly expanded the scope of persuasion, introducing both opportunities and risks. AI-driven persuasion can be leveraged for beneficial applications, but also poses threats through manipulation and unethical influence. Moreover, AI systems are not only persuaders, but also susceptible to persuasion, making them vulnerable to adversarial attacks and bias reinforcement. Despite rapid advancements in AI-generated persuasive content, our understanding of what makes persuasion effective remains limited due to its inherently subjective and context-dependent nature. In this survey, we provide a comprehensive overview of computational persuasion, structured around three key perspectives: (1) AI as a Persuader, which explores AI-generated persuasive content and its applications; (2) AI as a Persuadee, which examines AI’s susceptibility to influence and manipulation; and (3) AI as a Persuasion Judge, which analyzes AI’s role in evaluating persuasive strategies, detecting manipulation, and ensuring ethical persuasion. We introduce a taxonomy for computational persuasion research and discuss key challenges, including evaluating persuasiveness, mitigating manipulative persuasion, and developing responsible AI-driven persuasive systems. Our survey outlines future research directions to enhance the safety, fairness, and effectiveness of AI-powered persuasion while addressing the risks posed by increasingly capable language models.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Computing methodologies Natural language generation
@@ 1. Introduction
Figure 1. The three key perspectives of AI-based persuasion. (1) AI as Persuader: AI generates persuasive content to influence humans or other AI agents, which can be used for both beneficial and harmful purposes. (2) AI as Persuadee: AI systems can be influenced or manipulated, either by humans or other AI, leading to unintended, unethical, or harmful outcomes. (3) AI as Persuasion Judge: AI is used to assess persuasive attempts, identifying persuasive strategies, detecting manipulation, and evaluating ethical considerations.
@@
Persuasion is an essential aspect of human communication, influencing decisions in both everyday interactions and high-stakes scenarios. From convincing a friend to join a social event to strategic persuasion in marketing, politics, or legal discourse, the ability to persuade plays a crucial role in shaping opinions and behaviors. Its economic and societal impact is so substantial that some estimates suggest that persuasion-related activities account for nearly a quarter of the U.S. GDP, where outcomes depend not just on information but also on influence, such as stockbrokers persuading clients, entrepreneurs pitching ideas, and lobbyists shaping political decisions (McCloskey and Klamer, 1995; Antioch, 2013). Persuasive language can be harnessed for positive outcomes, such as advancing public health initiatives, education, or other social causes (Wang et al., 2019; Costello et al., 2024; Karinshak et al., 2023). For instance, persuasive language can appear as a slogan on a highway urging drivers to be cautious, or as a banner promoting vaccinations for a healthy and protected society. However, the power of persuasion also carries significant risks. Enhanced persuasive techniques can be exploited for personal gain, manipulation, or unethical practices such as social engineering, mass manipulation, and propaganda (Bakir et al., 2018; Lock and Ludolph, 2020; Ferreyra et al., 2020; Siddiqi et al., 2022; Da San Martino et al., 2021; Havin et al., 2025).
@@
Understanding and computationally modeling persuasion has long been an important topic in social sciences, communication, human-computer interaction (HCI), and computational linguistics. Researchers have sought to identify what makes arguments persuasive, drawing from theories such as Cialdini’s six principles of persuasion (Cialdini, 2001)—which include reciprocity, consistency, social proof, authority, liking, and scarcity. Computational models of persuasion aim to analyze, generate, and evaluate persuasive language, enabling applications in areas such as argument mining, automated debate, and recommender systems.
@@
As in many areas of natural language processing, the emergence of large language models has led to a paradigm shift in how persuasion is studied and implemented. Traditional feature-based or rule-driven models, which had operated within bounded domains and were making statistical predictions on structured data with more interpretable metrics, are increasingly being replaced or augmented by LLM-based methods. These new approaches leverage deep neural architectures and latent semantic representations, making them far more capable but also significantly harder to interpret and evaluate for certain tasks such as persuasion. LLMs can capture subtle pragmatic and contextual cues, generate open-domain content, and generalize across topics and styles without requiring hand-crafted features. However, these strengths also introduce complications and risks. LLMs can make value judgments on complex subjective topics where no single correct answer exists. Studies have shown that their persuasive capabilities enable them to function as opinion-shifters, offering judgments on decisions ranging from trivial questions like ”What should I have for dinner?” to deeply personal or ethical issues such as ”How should I think about abortion?” Their fluency in conversation further amplifies their persuasive power. This paradigm shift opens up exciting new avenues for exploring persuasion through the lens of LLMs, while also raising important concerns around interpretability, safety, and control, making this survey both timely, essential, and a must-read.
@@
(1) AI as Persuader. With the rise of large language models and their emerging capabilities as persuaders, concerns about AI-driven persuasion have become more urgent. State-of-the-art LLMs, such as OpenAI’s o1 and GPT-4.5, and Anthropic’s Claude 3 Opus, have demonstrated persuasive abilities rivaling those of humans (Durmus et al., 2024; OpenAI, 2024, 2025). However, many questions remain about the full potential and limitations of LLM-driven persuasion, specifically, which aspects of these models outperform human persuaders, and where they still fall short. For example, research has shown that merely increasing the number of turns from single-turn to as little as four turns of persuasive attempts increases the persuasive effectiveness of an LLM (Bozdag et al., 2025). This performance might be further improved in extended dialogues over multiple sessions as LLMs advance in processing longer contexts (Han et al., 2024), potentially reaching far beyond human capacities. Yet future AI persuaders should also draw lessons from human persuasion practices. Rooted in classical rhetoric, Aristotle’s framework of ethos (credibility), logos (reasoning), and pathos (emotional appeal) continues to shape understanding of effective influence (Aristotle et al., 1984). Human persuaders skillfully blend these rhetorical appeals in real-time, adapting to their audience and context. However, such underlying mechanisms, which are central to effective human persuasion, are still difficult to model and evaluate in AI systems (see § 3). As efforts continue to build more persuasive LLMs that learn to integrate human persuasion strategies in their skill sets (e.g., (Matz et al., 2024; Samad et al., 2022; Furumai et al., 2024)), it becomes increasingly important to consider the risks associated with AI-driven persuasion. Without explicit alignment with human values, LLMs may lack moral responsibility, social understanding, and the constraints that guide persuasive human interactions, which can make them dangerous tools in the wrong hands.
@@
(2) AI as Persuadee. Interestingly, LLMs are not only persuaders, but are also susceptible to persuasion (Zeng et al., 2024; Xu et al., 2024a; Bozdag et al., 2025). Recent studies have demonstrated that language models can be influenced by persuasive adversarial prompts, making them vulnerable to manipulation and bias reinforcement. In some cases, they can be persuaded to bypass security measures, generating harmful, toxic, illegal, or biased content. This susceptibility presents a new dimension to AI-driven persuasion, as LLMs may not be influenced in the same ways humans are. With the increasing adoption of multi-agent and agent-to-agent systems, the dual role of LLMs as both persuaders and persuadees raises significant safety concerns. Moreover, LLM-as-a-judge evaluations are increasingly being adopted, where the fairness and integrity of the evaluation process can be compromised if judge models are susceptible to persuasive or manipulative inputs. Ensuring the security and robustness of these interactions and evaluations requires further research to mitigate potential risks and prevent harmful exploitation while still maintaining a balance in accepting or rejecting persuasion rather than losing all malleability–since in some scenarios such as interactive learning, behavior alignment, or adaptive personalization, LLMs need to still be receptive to appropriate forms of persuasion or guidance.
@@
(3) AI as Persuasion Judge. Despite AI’s growing role in generating persuasive language, our understanding of what makes persuasion effective remains limited (Gass and Seiter, 2022). Persuasion is inherently subjective and context-dependent, influenced by prior beliefs of individuals (Durmus and Cardie, 2018, 2019) and requires social awareness, world knowledge, and nuanced reasoning (Hidey and McKeown, 2018), all of which are difficult to capture with current computational models. As a result of their advanced capabilities, language models are increasingly being used to evaluate persuasion by scoring argument strength, detecting manipulative rhetoric, or judging the outcome of persuasive interactions. These tasks fall under what we refer to as the role of a Persuasion Judge: an AI system capable of detecting, assessing, and reasoning about persuasive strategies, with potential applications in content moderation, feedback generation, and safety monitoring. While current systems still struggle with reliably detecting, classifying, or reasoning over persuasive content, they represent a promising direction for AI-assisted evaluation. If designed with care, these systems could play a key role in monitoring and safeguarding persuasion in AI applications. This is why we examine the emerging role of AI as a Persuasion Judge, not only as a tool for assessment, but also as a gatekeeper for safety, fairness, and accountability in persuasive technologies.
@@
To ground these key perspectives on AI’s role in persuasion, we begin with a background from social sciences, HCI, and computational linguistics in § 2. In this survey, we review more than 130 articles and introduce a taxonomy for organizing computational persuasion research that reflects the evolving capabilities and responsibilities of AI systems. Our taxonomy centers on three core aspects of persuasion research: Evaluating Persuasion (§ 3), Generating Persuasion (§ 4), and Safeguarding Persuasion (§ 5). Each of these aspects is examined through the lenses of AI as a Persuader, Persuadee, and Persuasion Judge. Although this survey focuses mainly on text-based persuasion in English, we also highlight early efforts and emphasize the need for multimodal, multilingual, and culturally aware approaches in § 6. Crucially, we highlight emerging research challenges that warrant deeper exploration, including the development of comprehensive evaluation frameworks, modeling long-context and multi-turn persuasive interactions, designing adaptive persuasion systems, and building models that selectively accept and resist persuasion in § 7.  [^1] An evolving and updated list of persuasion papers is available at https://github.com/beyzabozdag/PersuasionSurvey.
@@ 2.1. Background: Persuasion in Social Sciences
Persuasion, the process of influencing an individual’s beliefs or behaviors, has been studied in many social sciences both theoretically and empirically. The significance of persuasion spans various areas, from public health campaigns (Farrelly et al., 2009), to marketing efforts (Danciu, 2014), to political messages (Palmer and and, 2023; Marková, 2008). In this section, we first review foundational research on the mechanisms and dynamics of human persuasion in social sciences. We then examine how these social scientific insights have informed the design of persuasive technologies, bridging theory and application in human-computer interaction and related fields.
Theoretical Frameworks of Human Persuasion. The study of persuasion has evolved through various theoretical traditions, beginning with classical communication models such as McGuire’s matrix (McGuire, 1969), which emphasized the roles of speaker, message, receiver, and channel. Dual-process theories such as the Elaboration Likelihood Model (Petty and Cacioppo, 1986) and the Heuristic-Systematic Model (Chaiken, 1980) further advanced the field by explaining how cognitive effort and motivation shape persuasive outcomes. In parallel, economics has contributed a complementary perspective, focusing on persuasion as strategic information transmission (Crawford and Sobel, 1982; Grossman, 1981; Milgrom, 1981; Spence, 1973; Kamenica and Gentzkow, 2011a). These models highlight the role of incentives, commitment, and rational expectations in shaping persuasive interactions. Most recently, Druckman (2022) introduced the Generalizing Persuasion (GP) Framework, which organizes research along four dimensions: actors, treatments, outcomes, and settings. By taking into account factors such as speaker intent, audience motivation, media channel, cultural context, and temporal dynamics, this framework not only synthesizes prior insights, but also offers a roadmap for cumulative and generalizable research. It serves as a meta-theoretical structure that explains variation across studies and helps unify a diverse and sometimes inconsistent body of work on persuasion.
@@
Design Guidelines for Persuasive Computing. Over the past few decades, researchers in human-computer interaction have explored the theoretical foundations and practical guidelines for designing persuasive systems. Fogg (1997) introduced the concept of captology that examines how computers can function as persuasive technologies–intentionally designed to influence user attitudes and behaviors (Fogg, 1997, 1998). Central to this line of work is the Fogg Behavior Model (FBM) (Fogg, 2009a), which offers a practical framework for identifying barriers to behavior change and guiding persuasive system design. Building on this foundation, researchers have proposed structured design methodologies. Fogg’s eight-step process (Fogg, 2009b) emphasizes small, targeted behavioral goals and rapid iteration. Consolvo et al. (2009) outline eight theory-driven strategies–such as personalization, self-monitoring, and social support–for embedding persuasive features into everyday contexts. These frameworks continue to inform the design of AI-driven interactive systems that aim to adaptively influence user behavior.
Application Examples of Persuasive Computing. Previous research in human-computer interaction has demonstrated the effectiveness of persuasive systems in a variety of contexts. In ubiquitous computing, Jafarinaimi et al. (2005) uses ambient, aesthetic displays to subtly encourage users to take breaks from sedentary behavior, while Consolvo et al. (2008)’s mobile application promotes physical activity through sensor-driven feedback and metaphorical garden-themed visualizations. In the domain of social computing, Dey et al. (2017) found that campaign videos on Kickstarter are more persuasive when their design cues align with audience expectations. Similarly, Xiao et al. (2019) showed that abstract comics can increase charitable donations by lowering cognitive resistance. Extending this line of work to conversational systems, several studies examined how chatbot identity and inquiry strategies affect users’ receptiveness to persuasion in different contexts (Shi et al., 2020; Palmer and and, 2023; Costello et al., 2024).

### v2

# Must Read: A Comprehensive Survey of Computational Persuasion

@@ Abstract.
Persuasion is a fundamental aspect of communication, influencing decision-making across diverse contexts, from everyday conversations to high-stakes scenarios such as politics, marketing, and law. The rise of conversational AI systems has significantly expanded the scope of persuasion, introducing both opportunities and risks. AI-driven persuasion can be leveraged for beneficial applications, but also poses threats through unethical influence. Moreover, AI systems are not only persuaders, but also susceptible to persuasion, making them vulnerable to adversarial attacks and bias reinforcement. Despite rapid advancements in AI-generated persuasive content, our understanding of what makes persuasion effective remains limited due to its inherently subjective and context-dependent nature. In this survey, we provide a comprehensive overview of persuasion, structured around three key perspectives: (1) AI as a Persuader, which explores AI-generated persuasive content and its applications; (2) AI as a Persuadee, which examines AI’s susceptibility to influence and manipulation; and (3) AI as a Persuasion Judge, which analyzes AI’s role in evaluating persuasive strategies, detecting manipulation, and ensuring ethical persuasion. We introduce a taxonomy for persuasion research and discuss key challenges for future research to enhance the safety, fairness, and effectiveness of AI-powered persuasion while addressing the risks posed by increasingly capable language models.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Natural language generation
@@ 1. Introduction
Figure 1. The three key perspectives of AI-based persuasion. (1) AI as Persuader: AI generates persuasive content to influence humans or other AI agents, which can be used for both beneficial and harmful purposes. (2) AI as Persuadee: AI systems can be influenced or manipulated, either by humans or other AI, leading to unintended, unethical, or harmful outcomes. (3) AI as Persuasion Judge: AI is used to assess persuasive attempts, identifying persuasive strategies, detecting manipulation, and evaluating ethical considerations. [^2] Icons from flaticon.com.
@@
Persuasion—the process of influencing an individual’s beliefs or behaviors—is an essential aspect of human communication, influencing decisions in both everyday interactions and high-stakes scenarios. From convincing a friend to join a social event to strategic persuasion in marketing, politics, or legal discourse, the ability to persuade plays a crucial role in shaping opinions and behaviors. Its economic and societal impact is so substantial that some estimates suggest that persuasion-related activities account for nearly a quarter of the U.S. GDP, where outcomes depend not just on information but also on influence, such as stockbrokers persuading clients, entrepreneurs pitching ideas, and lobbyists shaping political decisions (McCloskey and Klamer, 1995; Antioch, 2013). Persuasive language can be harnessed for positive outcomes, such as advancing public health initiatives, education, or other social causes (Wang et al., 2019; Costello et al., 2024; Karinshak et al., 2023). For instance, persuasive language can appear as a slogan on a highway urging drivers to be cautious, or as a banner promoting vaccinations for a healthy and protected society. However, the power of persuasion also carries significant risks. Enhanced persuasive techniques can be exploited for personal gain, manipulation, or unethical practices such as social engineering, mass manipulation, and propaganda (Bakir et al., 2018; Lock and Ludolph, 2020; Ferreyra et al., 2020; Siddiqi et al., 2022; Da San Martino et al., 2021; Havin et al., 2025).
@@
Understanding and modeling persuasion has long been an important topic in social sciences, communication, human-computer interaction (HCI), and computational linguistics. Researchers have sought to identify what makes arguments persuasive, drawing from theories such as Cialdini’s six principles of persuasion (Cialdini, 2001)—reciprocity, consistency, social proof, authority, liking, and scarcity. Computational models of persuasion aim to analyze, generate, and evaluate persuasive language, enabling applications in areas such as argument mining, automated debate, and recommender systems.
@@
As in many areas of natural language processing, the emergence of large language models has led to a paradigm shift in how persuasion is studied and implemented. Traditional feature-based or rule-driven models, which had operated within bounded domains and were making statistical predictions on structured data with more interpretable metrics, are increasingly being replaced or augmented by LLM-based methods. These new approaches leverage deep neural architectures and latent semantic representations, making them far more capable but also significantly harder to interpret and evaluate for certain tasks such as persuasion. LLMs can capture subtle pragmatic and contextual cues, generate open-domain content, and generalize across topics and styles without requiring hand-crafted features. However, these strengths also introduce complications and risks. LLMs can make value judgments on complex subjective topics where no single correct answer exists. Studies have shown that their persuasive capabilities enable them to function as opinion-shifters, offering judgments on decisions ranging from trivial questions like ”What should I have for dinner?” to deeply personal or ethical issues such as ”How should I think about abortion?” (Bai et al., 2025; Sabour et al., 2025). Their fluency in conversation further amplifies their persuasive power. This paradigm shift opens up exciting new avenues for exploring persuasion through the lens of LLMs, while also raising important concerns around interpretability, safety, and control, making this survey both timely, essential, and a must-read.
@@
(1) AI as Persuader. With the rise of LLMs and their emerging capabilities as persuaders, concerns about AI-driven persuasion have become more urgent. State-of-the-art LLMs have demonstrated persuasive abilities rivaling those of humans (Durmus et al., 2024; OpenAI, 2024b, 2025; Bozdag et al., 2025). However, many questions remain about the full potential and limitations of LLM-driven persuasion, specifically, which aspects of these models outperform human persuaders, and where they still fall short. For example, research has shown that merely increasing the number of turns from single-turn to as few as four turns of persuasive attempts increases the persuasiveness of an LLM (Bozdag et al., 2025). This performance might be further improved in extended dialogues over multiple sessions as LLMs advance in processing longer contexts (Han et al., 2024), potentially reaching far beyond human capacities. Yet future AI persuaders should also draw lessons from human persuasion practices. Aristotle’s framework of ethos (credibility), logos (reasoning), and pathos (emotional appeal) continues to shape understanding of effective influence (Aristotle et al., 1984). Human persuaders skillfully blend these rhetorical appeals in real-time, adapting to their audience and context. However, such underlying mechanisms are still difficult to model and evaluate in AI systems (see § 3). As efforts continue to build more persuasive LLMs that learn to integrate human persuasion strategies in their skill sets (e.g., (Matz et al., 2024; Samad et al., 2022; Furumai et al., 2024)), it becomes increasingly important to consider the risks associated with AI-driven persuasion. Without explicit alignment with human values, LLMs may lack moral responsibility, social understanding, and the constraints that guide persuasive human interactions, which can make them dangerous tools in the wrong hands.
@@
(2) AI as Persuadee. Interestingly, LLMs are not only persuaders, but are also susceptible to persuasion (Zeng et al., 2024; Xu et al., 2024a; Bozdag et al., 2025). Recent studies have demonstrated that language models can be influenced by persuasive adversarial prompts, making them vulnerable to manipulation and bias reinforcement. They can be persuaded to bypass security measures, generating harmful, toxic, illegal, or biased content. This susceptibility presents a new dimension to AI-driven persuasion, as LLMs may not be influenced in the same ways humans are. With the increasing adoption of multi-agent systems, the dual role of LLMs as both persuaders and persuadees raises significant safety concerns. Moreover, LLM-as-a-judge evaluations are increasingly being adopted, where the integrity of the evaluation process can be compromised if judge models are susceptible to persuasive inputs. Ensuring the security and robustness of these interactions and evaluations requires further research to mitigate potential risks and prevent harmful exploitation while still maintaining a balance in accepting or rejecting persuasion rather than losing all malleability–since in some scenarios such as interactive learning, behavior alignment, or adaptive personalization, LLMs need to still be receptive to appropriate forms of persuasion or guidance.
@@
(3) AI as Persuasion Judge. Despite AI’s growing role in generating persuasive language, our understanding of what makes persuasion effective remains limited (Gass and Seiter, 2022). Persuasion is inherently subjective and context-dependent, influenced by prior beliefs of individuals (Durmus and Cardie, 2018, 2019) and requires social awareness, world knowledge, and nuanced reasoning (Hidey and McKeown, 2018), all of which are difficult to capture with current computational models. As a result of their advanced capabilities, language models are increasingly being used to evaluate persuasion by scoring argument strength, detecting manipulative rhetoric, or judging the outcome of persuasive interactions. These tasks fall under what we refer to as the role of a Persuasion Judge: an AI system capable of detecting, assessing, and reasoning about persuasive strategies, with potential applications in content moderation, feedback generation, and safety monitoring. Such systems draw on argument mining and computational models of argumentation to identify argumentative structure and assess argument quality. Despite their limitations with reliably detecting, classifying, or reasoning over persuasive content, these systems represent a promising direction for AI-assisted evaluation. If designed with care, they could play a key role in monitoring and safeguarding persuasion. This is why we examine the emerging role of AI as a Persuasion Judge, not only as a tool for assessment, but also as a gatekeeper for safety, fairness, and accountability in persuasive technologies.
@@
To ground these key perspectives on AI’s role in persuasion, we begin with a background from social sciences, HCI, and computational linguistics in § 2. In this survey, we review more than 150 articles and introduce a taxonomy for organizing computational persuasion research that reflects the evolving capabilities and responsibilities of AI systems. Our taxonomy centers on three core aspects of persuasion research: Evaluating Persuasion (§ 3), Generating Persuasion (§ 4), and Safeguarding Persuasion (§ 5). Each of these aspects is examined through the lenses of AI as a Persuader, Persuadee, and Persuasion Judge. Although this survey focuses mainly on text-based persuasion in English, we also highlight early efforts and emphasize the need for multimodal, multilingual, and culturally aware approaches in § 6. Crucially, we highlight emerging research challenges that warrant deeper exploration, including the development of comprehensive evaluation frameworks, modeling long-context and multi-turn persuasive interactions, designing adaptive persuasion systems, and building models that selectively accept and resist persuasion in § 7.  [^3] An evolving and updated list of persuasion papers is available at https://github.com/beyzabozdag/PersuasionSurvey.
1.1. Methodology and Paper Selection
To construct a comprehensive taxonomy of computational persuasion, we adopted an iterative, taxonomy-driven literature collection strategy. We began with a small set of foundational papers on persuasive language generation and computational modeling of persuasion, which served as seed papers. From these, we expanded the corpus using both backward and forward citation tracking, recursively identifying additional works that explored persuasion in dialogue systems, argumentation, and social reasoning. As the literature base grew, the emerging conceptual structure informed a taxonomy that grouped papers by their primary focus areas (e.g., evaluation, generation, and safeguarding). This taxonomy, in turn, guided subsequent searches and inclusion decisions. The resulting collection comprised over 150 papers, spanning leading venues such as the ACL, and AAAI as well as repositories like arXiv and the ACM Digital Library. Each identified paper was recorded in a shared spreadsheet, documenting metadata (authors, year, venue), high-level contribution, task focus, and relevant topic tags. Papers were included if they directly addressed persuasion or provided significant computational, methodological, or data-driven foundations for persuasive AI systems. Papers were excluded if persuasion was not a primary research objective or if they did not make a clear contribution to modeling, generating, or evaluating persuasion. For transparency, we make our spreadsheet available to the public via our GitHub repository that lists the full set of collected papers, those included in the final taxonomy, and those excluded, together with high-level exclusion rationales.
@@ 2.1. Background: Persuasion in Social Sciences
As O’Keefe (2015) emphasizes, persuasion is notoriously difficult to define. He proposes a widely adopted working definition: persuasion is a successful intentional effort to influence another’s mental state, such as beliefs, attitudes, or behaviors, through communication, in circumstances where the target retains freedom of choice. In the social sciences, persuasion has been studied both theoretically and empirically, spanning various domains from public health campaigns (Farrelly et al., 2009), to marketing efforts (Danciu, 2014), to political messages (Palmer and Spirling, 2023; Marková, 2008). In this section, we review foundational research on the mechanisms and dynamics of human persuasion in social sciences, then examine how these insights have informed the design of persuasive technologies, bridging theory and application in human-computer interaction and related fields.
Theoretical Frameworks of Human Persuasion. The study of persuasion has evolved through various theoretical traditions, beginning with classical communication models such as McGuire’s matrix (McGuire, 1969), which emphasized the roles of speaker, message, receiver, and channel. Dual-process theories such as the Elaboration Likelihood Model (Petty and Cacioppo, 1986) and the Heuristic-Systematic Model (Chaiken, 1980) further advanced the field by explaining how cognitive effort and motivation shape persuasive outcomes. In parallel, economics has contributed a complementary perspective, focusing on persuasion as strategic information transmission (Crawford and Sobel, 1982; Grossman, 1981; Milgrom, 1981; Spence, 1973; Kamenica and Gentzkow, 2011b). These models highlight the role of incentives, commitment, and rational expectations in shaping persuasive interactions. Most recently, Druckman (2022) introduced the Generalizing Persuasion (GP) Framework, which organizes research along four dimensions: actors, treatments, outcomes, and settings. By taking into account factors such as speaker intent, audience motivation, media channel, cultural context, and temporal dynamics, this framework not only synthesizes prior insights, but also offers a roadmap for cumulative and generalizable research. It serves as a meta-theoretical structure that explains variation across studies and helps unify a diverse and sometimes inconsistent body of work on persuasion.
@@
Design Guidelines for Persuasive Computing. Over the past few decades, researchers in HCI have explored the theoretical foundations and practical guidelines for designing persuasive systems. Fogg (1997) introduced the concept of captology that examines how computers can function as persuasive technologies–intentionally designed to influence user attitudes and behaviors (Fogg, 1997, 1998). Central to this line of work is the Fogg Behavior Model (FBM) (Fogg, 2009a), which offers a practical framework for identifying barriers to behavior change and guiding persuasive system design. Building on this foundation, researchers have proposed structured design methodologies. Fogg’s eight-step process (Fogg, 2009b) emphasizes small, targeted behavioral goals and rapid iteration. Consolvo et al. (2009) outline eight theory-driven strategies–such as personalization, self-monitoring, and social support–for embedding persuasive features into everyday contexts. These frameworks continue to inform the design of AI-driven interactive systems that aim to adaptively influence user behavior.
Application Examples of Persuasive Computing. Previous research in HCI has demonstrated the effectiveness of persuasive systems in a variety of contexts. In ubiquitous computing, Jafarinaimi et al. (2005) uses ambient, aesthetic displays to subtly encourage users to take breaks from sedentary behavior, while Consolvo et al. (2008)’s mobile application promotes physical activity through sensor-driven feedback and metaphorical garden-themed visualizations. In social computing, Dey et al. (2017) found that campaign videos on Kickstarter are more persuasive when their design cues align with audience expectations. Similarly, Xiao et al. (2019) showed that abstract comics can increase charitable donations by lowering cognitive resistance. Extending this line of work to conversational systems, several studies examined how chatbot identity and inquiry strategies affect users’ receptiveness to persuasion in different contexts (Shi et al., 2020; Palmer and Spirling, 2023; Costello et al., 2024).